How many L of a 65% sugar solution must be mixed with 9 L of a 85% sugar solution to make a 77% solution?
Answer by macston(5194)

.
S=amount of 65%; E=amount of 85%
.
0.65S+0.85E=0.77(S+E)
0.65S+0.85(9)=0.77(S+9)
0.65S+7.65=0.77S+6.93
0.72=0.12S
6=S
.
ANSWER: 6 liters of 65% solution should be used.
.